# Scalable Physics-Informed Neural Networks for Accelerating Electromagnetic Transient Stability Assessment

### v1

@@ Abstract
Submitted to the 24th Power Systems Computation Conference (PSCC 2026). This work was supported by the ERC Starting Grant VeriPhIED, Grant Agreement 949899, and the ERC Proof of Concept PINNSim, Grant Agreement 101248667, both funded by the European Research Council.
@@ I Introduction
The rapid expansion of Inverter-Based Resources (IBRs) in modern power systems, such as wind farms, photovoltaic systems, or batteries, has led to a significant increase in the computational requirements and costs associated with stability studies. For decades, power systems have relied on synchronous generators. IBRs pose a complete paradigm shift, based on a different technology that purely relies on control logic and high-frequency switching semiconductors. This shift presents a fundamental challenge to the power system industry, as fast-reacting and complex IBR dynamics replace the slow electromechanical ones [stabilityupdate, milanostability, stott]. Thus, time-domain simulations, the core method for stability studies, are caught between two opposing forces. On the one side, grid operators require more simulations and scenarios to ensure system reliability. On the other side, the cost of the required simulations drastically increases, as they feature more detailed and sophisticated models that often need Electromagnetic Transient (EMT) modeling.
While a few ideas have been proposed to address the rising cost of the much-needed time-domain simulations [emtaccelerationreview], the core of the simulation algorithms still remains untouched. The most applied approach, so far, has been using more powerful and dedicated hardware, which also opened the door to real-time simulations [rtparallel, rtds] and the parallelization of some processes [paraemt, parallelemt]. However, the primary methods used to simulate are the same.
New methods are needed to improve simulation performance. Physics-Informed Neural Networks (PINNs) have recently emerged as a powerful, fast, and flexible computational paradigm for capturing complex physical systems with high fidelity [raissi, neuralnetsdyn]. By embedding physical laws into the learning process, PINNs can accurately represent nonlinear system behaviors over extended time periods. These features make PINNs highly attractive for power system simulations as a new computing method that can significantly accelerate time-domain simulations.
Several works have addressed how to leverage PINNs to accelerate power system Root Mean Square (RMS) simulations, often known as phasor-type simulations [pinnsim, plugandplay, daepinn, powertechnew]. However, their applicability to EMT studies has remained an open question. The challenges associated with the highly nonlinear, sophisticated, and oscillatory nature of EMT simulations make PINN modeling very complex. This paper introduces a new methodology that enables a modular and efficient integration of PINNs for faster EMT simulations. PINNs, trained on the underlying physical equations, capture the closed-form solution of the most computationally intensive control logic, significantly reducing the computational burden at every time step. The contributions of this paper are as follows:
@@
The code is made available online to foster the adoption of PINNs in the power systems community [github_PINN].
@@ II-A The vision
PINNs have several characteristics that make them highly attractive to power system simulations. They can accurately learn nonlinear dynamics over large time steps and provide a fast, explicit, direct-solution model, thereby avoiding traditional stepping algorithms. Leveraging these characteristics is still an open research question for the power systems industry. One approach is to develop a new data-driven solver entirely based on PINNs. This was first explored in [pinnsim], showing significant speed-ups but requiring widespread PINN adoption. A second approach is to modularly integrate these speed-ups into the existing simulation framework [plugandplay]. This approach can still achieve significant speed-ups while remaining compatible with existing tools, and is therefore widely applicable.
@@
We classify power system simulations into two groups: RMS simulations and EMT simulations. This classification, similar to [revisiting], groups simulations by modeling detail. RMS simulations assume steady-state operation of the network, using phasors to capture electromechanical deviations. By contrast, EMT simulations model the waveforms of all phases across the system, requiring more detailed and sophisticated models. Therefore, EMT simulations are significantly more computationally expensive but capture dynamics over a broad band of frequencies. Despite fundamental differences between these two groups, the simulation speed of both algorithms mostly comes down to two factors: (i) how many time steps are required, and (ii) how much time it takes to evaluate all the system models in each time step.
In RMS simulations, the system models are relatively simple, as they typically use a positive-sequence single-phase equivalent. Thus, the dominant time cost comes from the number of time steps required, which is determined by how large the time steps are. Traditional methods quickly become inaccurate when the step size increases, limiting the solver’s step sizes and ultimately the simulation speed. PINNs, on the other hand, once trained with the underlying component dynamics, provide accurate representations for large steps. Considering that the fast-evolving components are the ones that constrain the maximum accurate step size, we can train PINNs for these fast components and achieve larger simulation step sizes without losing accuracy, as shown in [powertechnew].
@@ III-A Generic formulation
where $\textbf{x}(t)$ represents the differential states of the network equations, and $\textbf{y}(t)$ represents the differential states of the device model equations. $\textbf{u}(t)$ represents the algebraic states of the device models. Array m captures the parameters present in the network equations, and n in the device equations [ivp_daes].
@@ III-C EMT simulation program
There are several approaches to solve an EMT simulation problem. In this work, we consider the partitioned algorithm, used in most state-of-the-art power system software. This algorithm divides the EMT simulation problem into two parts, the electric network and the control systems [emtpbook].
@@ III-C1 Electrical systems
The electrical system solution is based on the principles outlined by H.W. Dommel in [dommel, emtpbookog]. The differential equations of the circuit elements in (2) are transformed into linearized algebraic equations by discretizing them with the trapezoidal rule. The resulting equations define the so-called companion models and depend on voltage, current, and past values. Fig. 2 depicts the companion models of an inductance and a capacitance. See [emtpbook] for more details on the transformation.
@@
where $\textbf{v}(t)$ are the bus voltages, $\textbf{i}(t)$ the branch currents, and $\textbf{I}(t)$ the history currents. Y is the symmetric nodal admittance matrix. Nonlinear electrical elements, such as transformers or surge arresters, can also be included in this formulation [emtpbook].
@@ III-C2 Control systems
The control systems of all devices in the system, usually defined by block diagrams, are solved together. These primarily include converters, excitation systems, and protective relays. The control systems solution algorithm, first proposed in [initacs] and labeled as TACS, also uses the trapezoidal rule to discretize its differential equations. How the control systems are solved then depends on the type of resulting system. The three possible types are solved as follows: (i) linear systems are solved analytically, (ii) systems with nonlinearities only in open-loop paths are solved sequentially, and (iii) systems where nonlinearities appear in closed loops require an iterative method or the introduction of a time delay to break the feedback loop.
Most EMT models result in the latter type, where nonlinearities appear in closed loops. The fastest and most common approach to solving such systems is to introduce a time delay in the feedback, thereby opening the loop to allow for a non-iterative sequential solution. The main shortcoming of this approach is reduced accuracy, and it can lead to instability in otherwise physically stable simulations [araujo, simultaneous2]. The most accurate approach is to use an iterative method to solve the closed-loop system; however, this approach has a higher computational cost. Fig. 3 illustrates both approaches.
@@ III-C3 Interface network-control solutions
Figure 4: Interface overview between the network and control systems solution [emtpbookog].
@@ IV-A General formulation
To parametrize a PINN, we use a fully connected feedforward Neural Network (NN) with $K$ hidden layers and $N_{k}$ neurons in each layer [Goodfellow]. Each hidden layer $k$ comprises the weight matrices $W^{k}$ and bias vectors $b^{k}$. Together, they define the set of learnable parameters of the PINN, $\Theta=[{W^{k},b^{k}}]$. The relation between inputs and outputs is defined by
@@ IV-E Training procedure
To obtain accurate PINN models, we optimize the training loss function for all weights and biases of the neural network. We utilize a loss function with data-based and physics-based components [raissi]. The data-based loss, $\mathcal{L}_{u}$, minimizes the difference between the target values and the PINN predictions using a mean squared error. These targets are provided by a dataset, $\mathcal{D}_{u}$, which contains accurate simulated results for different time step sizes and initial conditions:
@@ V-A Study case
We consider an aggregated type-4 wind turbine model connected to an external grid through an LCL filter [kazem]. We model the grid-side converter as an average controlled voltage source, assuming a constant DC voltage and neglecting the turbine generator and the rotor-side converter. Fig. 6 depicts the control architecture of the considered wind turbine model. A synchronous reference-frame PLL is implemented to synchronize with the external grid. We apply vector control in the dq frame, utilizing a cascaded PI configuration that controls, in the outer loop, the active and reactive power injected into the external grid. We consider an ideal modulation block. The system configuration and parameters are based on the CIGRE benchmark model C4.49 [Lukash]. Table I presents the parameters considered for the wind turbine system.
@@ V-B Study case implementation
We implement the PINN training and integration using the NumPy and Torch libraries [numpy, torch]. The trained PINN consists of 64 neurons in each of three hidden layers, with the tanh activation function. We trained the PINN for $10^{6}$ epochs with the Adam optimizer using a decaying learning rate. Training was performed on an NVIDIA Tesla A100 GPU [DTU_DCC_resource]. The setup parameters were determined empirically, iteratively tuning the architecture and the training algorithm to achieve the most accurate results on a testing dataset. We train the PINN for a broad band of operational conditions, enabling the PINN to be reused across all simulations and scenarios in the subsequent analysis.

### v2

@@ Abstract
This work was supported by the ERC Starting Grant VeriPhIED, Grant Agreement 949899, and the ERC Proof of Concept PINNSim, Grant Agreement 101248667, both funded by the European Research Council.
@@ I Introduction
The rapid expansion of Inverter-Based Resources (IBRs) in modern power systems, such as wind farms, photovoltaic systems, or batteries, has led to a significant increase in the computational requirements and costs associated with stability studies. For decades, power systems have relied on synchronous generators. IBRs pose a complete paradigm shift, based on a different technology that purely relies on control logic and high-frequency switching semiconductors. This shift presents a fundamental challenge to the power system industry, as fast-reacting and complex IBR dynamics replace the slow electromechanical ones [12, 18, 24]. Thus, time-domain simulations, the core method for stability studies, are caught between two opposing forces. On the one side, grid operators require more simulations and scenarios to ensure system reliability. On the other side, the cost of the required simulations drastically increases, as they feature more detailed and sophisticated models that often need Electromagnetic Transient (EMT) modeling.
While a few ideas have been proposed to address the rising cost of the much-needed time-domain simulations [25], the core of the simulation algorithms still remains untouched. The most applied approach, so far, has been using more powerful and dedicated hardware, which also opened the door to real-time simulations [4, 20] and the parallelization of some processes [29, 5]. However, the primary methods used to simulate are the same.
New methods are needed to improve simulation performance. Physics-Informed Neural Networks (PINNs) have recently emerged as a powerful, fast, and flexible computational paradigm for capturing complex physical systems with high fidelity [22, 15]. By embedding physical laws into the learning process, PINNs can accurately represent nonlinear system behaviors over extended time periods. These features make PINNs highly attractive for power system simulations as a new computing method that can significantly accelerate time-domain simulations.
Several works have addressed how to leverage PINNs to accelerate power system Root Mean Square (RMS) simulations, often known as phasor-type simulations [23, 28, 19, 27]. However, their applicability to EMT studies has remained an open question. The challenges associated with the highly nonlinear, sophisticated, and oscillatory nature of EMT simulations make PINN modeling very complex. This paper introduces a new methodology that enables a modular and efficient integration of PINNs for faster EMT simulations. PINNs, trained on the underlying physical equations, capture the closed-form solution of the most computationally intensive control logic, significantly reducing the computational burden at every time step. The contributions of this paper are as follows:
@@
The code is made available online to foster the adoption of PINNs in the power systems community [26].
@@ II-A The vision
PINNs have several characteristics that make them highly attractive to power system simulations. They can accurately learn nonlinear dynamics over large time steps and provide a fast, explicit, direct-solution model, thereby avoiding traditional stepping algorithms. Leveraging these characteristics is still an open research question for the power systems industry. One approach is to develop a new data-driven solver entirely based on PINNs. This was first explored in [23], showing significant speed-ups but requiring widespread PINN adoption. A second approach is to modularly integrate these speed-ups into the existing simulation framework [28]. This approach can still achieve significant speed-ups while remaining compatible with existing tools, and is therefore widely applicable.
@@
We classify power system simulations into two groups: RMS simulations and EMT simulations. This classification, similar to [14], groups simulations by modeling detail. RMS simulations assume steady-state operation of the network, using phasors to capture electromechanical deviations. By contrast, EMT simulations model the waveforms of all phases across the system, requiring more detailed and sophisticated models. Therefore, EMT simulations are significantly more computationally expensive but capture dynamics over a broad band of frequencies. Despite fundamental differences between these two groups, the simulation speed of both algorithms mostly comes down to two factors: (i) how many time steps are required, and (ii) how much time it takes to evaluate all the system models in each time step.
In RMS simulations, the system models are relatively simple, as they typically use a positive-sequence single-phase equivalent. Thus, the dominant time cost comes from the number of time steps required, which is determined by how large the time steps are. Traditional methods quickly become inaccurate when the step size increases, limiting the solver’s step sizes and ultimately the simulation speed. PINNs, on the other hand, once trained with the underlying component dynamics, provide accurate representations for large steps. Considering that the fast-evolving components are the ones that constrain the maximum accurate step size, we can train PINNs for these fast components and achieve larger simulation step sizes without losing accuracy, as shown in [27].
@@ III-A Generic formulation
where $\textbf{x}(t)$ represents the differential states of the network equations, and $\textbf{y}(t)$ represents the differential states of the device model equations. $\textbf{u}(t)$ represents the algebraic states of the device models. Array m captures the parameters present in the network equations, and n in the device equations [3].
@@ III-C EMT simulation program
There are several approaches to solve an EMT simulation problem. In this work, we consider the partitioned algorithm, used in most state-of-the-art power system software. This algorithm divides the EMT simulation problem into two parts, the electric network and the control systems [17].
@@ III-C1 Electrical systems
The electrical system solution is based on the principles outlined by H.W. Dommel in [7, 6]. The differential equations of the circuit elements in (2) are transformed into linearized algebraic equations by discretizing them with the trapezoidal rule. The resulting equations define the so-called companion models and depend on voltage, current, and past values. Fig. 2 depicts the companion models of an inductance and a capacitance. See [17] for more details on the transformation.
@@
where $\textbf{v}(t)$ are the bus voltages, $\textbf{i}(t)$ the branch currents, and $\textbf{I}(t)$ the history currents. Y is the symmetric nodal admittance matrix. Nonlinear electrical elements, such as transformers or surge arresters, can also be included in this formulation [17].
@@ III-C2 Control systems
The control systems of all devices in the system, usually defined by block diagrams, are solved together. These primarily include converters, excitation systems, and protective relays. The control systems solution algorithm, first proposed in [9] and labeled as TACS, also uses the trapezoidal rule to discretize its differential equations. How the control systems are solved then depends on the type of resulting system. The three possible types are solved as follows: (i) linear systems are solved analytically, (ii) systems with nonlinearities only in open-loop paths are solved sequentially, and (iii) systems where nonlinearities appear in closed loops require an iterative method or the introduction of a time delay to break the feedback loop.
Most EMT models result in the latter type, where nonlinearities appear in closed loops. The fastest and most common approach to solving such systems is to introduce a time delay in the feedback, thereby opening the loop to allow for a non-iterative sequential solution. The main shortcoming of this approach is reduced accuracy, and it can lead to instability in otherwise physically stable simulations [1, 16]. The most accurate approach is to use an iterative method to solve the closed-loop system; however, this approach has a higher computational cost. Fig. 3 illustrates both approaches.
@@ III-C3 Interface network-control solutions
Figure 4: Interface overview between the network and control systems solution [6].
@@ IV-A General formulation
To parametrize a PINN, we use a fully connected feedforward Neural Network (NN) with $K$ hidden layers and $N_{k}$ neurons in each layer [10]. Each hidden layer $k$ comprises the weight matrices $W^{k}$ and bias vectors $b^{k}$. Together, they define the set of learnable parameters of the PINN, $\Theta=[{W^{k},b^{k}}]$. The relation between inputs and outputs is defined by
@@ IV-E Training procedure
To obtain accurate PINN models, we optimize the training loss function for all weights and biases of the neural network. We utilize a loss function with data-based and physics-based components [22]. The data-based loss, $\mathcal{L}_{u}$, minimizes the difference between the target values and the PINN predictions using a mean squared error. These targets are provided by a dataset, $\mathcal{D}_{u}$, which contains accurate simulated results for different time step sizes and initial conditions:
@@ V-A Study case
We consider an aggregated type-4 wind turbine model connected to an external grid through an LCL filter [2]. We model the grid-side converter as an average controlled voltage source, assuming a constant DC voltage and neglecting the turbine generator and the rotor-side converter. Fig. 6 depicts the control architecture of the considered wind turbine model. A synchronous reference-frame PLL is implemented to synchronize with the external grid. We apply vector control in the dq frame, utilizing a cascaded PI configuration that controls, in the outer loop, the active and reactive power injected into the external grid. We consider an ideal modulation block. The system configuration and parameters are based on the CIGRE benchmark model C4.49 [13]. Table I presents the parameters considered for the wind turbine system.
@@ V-B Study case implementation
We implement the PINN training and integration using the NumPy and Torch libraries [11, 21]. The trained PINN consists of 64 neurons in each of three hidden layers, with the tanh activation function. We trained the PINN for $10^{6}$ epochs with the Adam optimizer using a decaying learning rate. Training was performed on an NVIDIA Tesla A100 GPU [8]. The setup parameters were determined empirically, iteratively tuning the architecture and the training algorithm to achieve the most accurate results on a testing dataset. We train the PINN for a broad band of operational conditions, enabling the PINN to be reused across all simulations and scenarios in the subsequent analysis.